count ← 0
for(i=n; i>0; i←⌊i/2⌋)

for(j=0; j<i; j++) count = count + 1

What’s the Big-Theta of the running time?

indicate your time complexity.
Answer #1

values of i of outer loop are n, n/2, n/4, n/8, ...
    Explanation:
    -------------
    because i value is changing by i = i* through each iteration
    so, i changes to 1, 2, 4, ..., n/4, n/2, n
inner loop iterates i times.

to find total number of iterations, add all values of i
all, possible values of i are 1, 2, 4, ..., n/4, n/2, n
so, total number of iterations = n+n/2+n/4+n/8+...+4+2+1
= n(1+1/2+1/4+1/8+...)
Note: 1+1/2+1/4+1/8+... <= 2
so, n(1+1/2+1/4+1/8+...) <= 2n
hence, time complexity is ?(n)
Answer: ?(n)